OpenStudy (anonymous):

Fred is fencing his yard, which is shaped as a rhombus. He used 4x meters of fencing material for one side and (x + 9) meters for the adjacent side. What is the perimeter of the yard? Answer 12 meters 36 meters 48 meters 96 meters

OpenStudy (anonymous):

Perimeter = 2l + 2w 2(4x) + 2(x + 9)

OpenStudy (anonymous):

10x + 18 ??

OpenStudy (anonymous):

Yes now find x

OpenStudy (anonymous):

ok umm what do i do to find it

OpenStudy (anonymous):

You want to find the value of x. So you have to put it alone on one side of the equation: 10x + 18 = 0 -18 10x = -18 /10 x = -1.8 Now put in x into the original equation to find the perimeter;) 2(4x) + 2(x + 9) 2(4(-1.8)) + 2((-1.8) + 9

OpenStudy (anonymous):

Hold on a sec

OpenStudy (anonymous):

i got -93.6 that cant be it

OpenStudy (anonymous):

Yeah I was just checking and it didn't work:( Sorry! Um here is the same question by someone else on openstudy and the response:) I checked on other places and this seems to be right:) http://openstudy.com/study#/updates/4fb09d5ee4b059b524fcc13b

OpenStudy (anonymous):

I'm reallyyyy sorry!

OpenStudy (anonymous):

ur fine i promise no big deal atleast u cought yourself i aint one of that will go off on you for giving me the wrong answer. I look at it as in atleast you were kind enough to help me

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

anytime:)